# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

### v3

@@ Abstract
Vision-language models (VLMs) have made significant progress in reasoning within natural scenes, yet their potential in medical imaging remains largely underexplored. Medical reasoning tasks, which require robust image analysis and the generation of well-justified answers, present unique challenges due to the inherent complexity of medical images. To be clinically useful, model outputs must be accurate, and logically consistent, as errors or inconsistencies can have critical consequences in medical applications. To address these challenges, we propose Med-R1, a novel framework that investigates whether reinforcement learning (RL) can enhance the generalizability and trustworthiness of VLMs in medical reasoning. Building on the recently introduced DeepSeek strategy, we adopt Group Relative Policy Optimization (GRPO) for RL, which encourages models to explore reasoning paths guided by reward signals. Unlike supervised fine-tuning (SFT), which often overfits to training data and struggles with generalization, RL enables models to develop more robust and diverse reasoning capabilities. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. Compared to the base model, Qwen2-VL-2B, Med-R1 achieves a 29.94% improvement in average accuracy across these modalities and even outperforms Qwen2-VL-72B—a model with 36 times more parameters. To assess model’s generalization abilities, we further test on five different question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 also demonstrates superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% and Qwen2-VL-72B in question-type generalization accuracy. These results highlight that RL not only enhances medical reasoning capabilities but also enables parameter-efficient models to exceed the performance of significantly larger counterparts. By demonstrating strong cross-domain and cross-tasks performance, Med-R1 points toward a new direction for developing practical and generalizable medical VLMs.
@@ 1 Introduction
Vision-language models (VLMs) have achieved remarkable progress in reasoning over natural images (e.g., GPT-4o [16], Gemini-1.5 [33], and Qwen-VL [2]), demonstrating impressive capabilities in tasks such as visual question answering (VQA) and multimodal dialogue. These advancements have been driven by large-scale vision-language pretraining and supervised fine-tuning (SFT), which enables models to associate visual content with textual information. However, applying VLMs to medical imaging reasoning remains a significant challenge. Unlike natural images, medical data demands precise interpretation, requiring VLMs to produce not only answer outcomes but also intermediate reasoning paths that are consistent with clinical decision-making logic. For instance, diagnosing a lung nodule in a CT scan necessitates localizing lesions, analyzing morphological features, and integrating patient history—a reasoning chain that must be both accurate and interpretable to gain clinicians’ trust. Furthermore, the diversity of medical imaging modalities (e.g., CT, MRI, microscopy) and task types (e.g., diagnosis, lesion grading) poses stringent demands on model generalizability. These challenges raise a critical question: How can we equip VLMs with robust reasoning abilities that generalize across medical domains while maintaining medical reliability?
@@
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [6, 1]. While SFT has been widely adopted to adapt foundation models to medical imaging [24, 36, 18, 5, 22], it suffers from two fundamental issues that hinder medical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [24, 36] lack access to high-quality CoT data, leading to shallow and non-clinically grounded reasoning. These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks. This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are are indispensable requirements.
To address these challenges, we propose Med-R1, a framework leveraging Reinforcement Learning (RL) [29] to enhance both the generalizability and interpretability of medical VLMs. In contrast to SFT, which primarily aligns outputs with fixed supervision, RL promotes exploration of alternative reasoning strategies by leveraging reward-driven feedback, allowing models to construct internal logic without requiring explicit CoT annotations. This approach mitigates overfitting and shortcut learning, as demonstrated by Chu et al. [9], who showed that RL-trained vision-language models significantly outperform SFT in cross-domain generalization. Building on recent advancements in RL optimization, we adopt Group Relative Policy Optimization (GRPO) [30], a novel strategy that eliminates the need for complex value models required by traditional methods like Proximal Policy Optimization (PPO) [29]. GRPO incorporates rule-based rewards derived from medical guidelines (e.g., diagnostic criteria for tumors) and group-relative comparisons to stabilize policy updates. For example, rewards can be designed to prioritize reasoning steps that align with radiological decision trees, ensuring clinically plausible rationales. This dual mechanism not only reduces computational overhead but also enforces adherence to domain-specific constraints, making GRPO uniquely suited for medical reasoning tasks where interpretability and scalability are essential for deployment.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding medical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
As shown in Figure 1, to rigorously validate the effectiveness of Med-R1, we conduct a comprehensive evaluation spanning eight clinically critical imaging modalities: Computed Tomography (CT), Magnetic Resonance Imaging (MRI), Ultrasound, Dermoscopy, Fundus Photography, Optical Coherence Tomography (OCT), Microscopy Images, and X-ray Imaging. These modalities were selected to represent the spectrum of medical imaging challenges, from macroscopic anatomical visualization (CT/MRI) to cellular-level analysis (Microscopy) and functional assessments (OCT). Across these domains, Med-R1 (2B parameters) achieves an average accuracy of 69.91%, representing a 29.94% absolute improvement over its foundation model Qwen2-VL-2B. Notably (Table 2), Med-R1 even surpasses Qwen2-VL-72B—a 72B-parameter model with 36× larger scale, demonstrating that parameter efficiency need not compromise performance when leveraging RL-driven reasoning. Beyond evaluating performance across imaging modalities, we also examine Med-R1’s generalizability across different medical question types. We assess the model’s capability in five key question categories: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 exhibits superior cross-task generalization, outperforming Qwen2-VL-2B by 32.06% in question-type generalization accuracy, and even surpassing Qwen2-VL-72B in this aspect (Table 4). These findings emphasize the potential of RL-based training in improving both reasoning depth and task adaptability in medical VLMs.
@@
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized the contribution as follows:
@@
    Generalizable Medical VLM with Multi-Modal Explicit Reasoning: We propose Med-R1, the first vision-language model supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.), capable of generating step-by-step reasoning explanations without task-specific retraining. Unlike prior SFT-based models limited to single modalities and black-box answers, Med-R1 achieves cross-domain generalization through RL-driven exploration.
@@
    Reward Guided Reasoning via Constrained RL: We employ Group Relative Policy Optimization, a reinforcement learning method that removes the need for explicit reasoning annotations. GRPO combines rule-informed reward signals (e.g., alignment with diagnostic protocols) with group-relative optimization, allowing models to align output behavior with medically grounded reasoning from final-answer supervision alone.
@@
    Robust and Efficient Generalization: Med-R1 outperforms the base foundation model by 29.94% and SFT baselines by 15.84% in cross-modality performance. It further improves 32.06% over the foundation model and 11.25% over SFT baselines on cross-task generalization. Med-R1 also exceeds much larger models in both general-domain and medical-domain settings—including Qwen2-VL-72B and MedVInT-7B—demonstrating scalable reasoning and medical transferability beyond model size.
@@ 2 Related Works
General VLMs and Medical VLMs. VLMs have significantly advanced multimodal understanding by aligning visual and textual representations through large-scale pretraining. General-purpose VLMs, such as CLIP [28] and BLIP-2 [19], excel in natural image-text tasks but encounter challenges when applied to domain-specific medical reasoning. Recent efforts have sought to adapt these models to medical imaging through SFT on domain-specific datasets. For instance, LLaVA [23] introduces instruction tuning for general-purpose vision-language tasks, achieving strong performance on natural image benchmarks. Building on this, LLaVA-Med [18] fine-tunes LLaVA using radiology reports, demonstrating improved performance on medical VQA tasks. Similarly, Med-Flamingo [24] extends Flamingo by incorporating domain-specific pretraining for medical VQA. However, these SFT-based approaches often suffer from overfitting to narrow medical corpora, limiting their generalizability across diverse imaging modalities and medical tasks. Our work addresses this gap by introducing RL for generalization, enabling robust performance without modality-specific retraining.
Reinforcement Learning for Post-Training. Reinforcement learning [32, 8] has emerged as a powerful tool for aligning large language models (LLMs) with human preferences through reward signals. Approaches like RLHF [25] and RLAIF [4] leverage human or AI-generated feedback to refine model behavior. In the vision-language domain, RL has been applied to improve VQA accuracy [9] and reduce hallucination [30]. However, existing methods rely on complex reward models or extensive human annotation, limiting scalability in data-scarce domains like medicine. The Group Relative Policy Optimization addresses these limitations by incorporating rule-based rewards derived from medical guidelines and group-relative comparisons, enabling parameter-efficient adaptation to diverse medical tasks.
Medical Reasoning and Interpretability Recent work emphasizes the importance of interpretable reasoning in medical AI, with approaches like CoT prompting [35] and program-guided reasoning [6]. However, acquiring such annotations in the medical domain is particularly challenging due to the need for expert involvement [21], high annotation costs, and the complexity of medical reasoning. The RL approach enables emergent reasoning without explicit supervision, addressing this challenge. Concurrent to our work, MedVLM-R1 [26] explores reinforcement learning for radiology VQA using a GRPO-based approach. While promising, the study is conducted on a small-scale dataset (600 training samples) and does not evaluate generalization across tasks or modalities, which makes it challenging to assess the robustness of the proposed reasoning capability and its generalizability. Nonetheless, this work also reflects the growing interest in applying reinforcement learning to medical vision-language models.
Figure 2: Medical VQA examples of Med-R1 The left panel demonstrates a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray). The right panel illustrates a disease diagnosis task, where Med-R1 accurately detects urolithiasis in an abdominal CT scan. In both cases, the model provides interpretable reasoning (“Think”), explaining its decision-making process before selecting the final answer.
@@ 3 Method
We adopt RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [15] and visual navigation [9]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [31] to train a large base MLLM [3] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the algorithm, our reward design, and data structure below.
@@ 3.1 Group Relative Policy Optimization (GRPO)
RL-based algorithms such as PPO [29] and GRPO [30] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities. Unlike supervised fine-tuning, which directly optimizes maximum likelihood, these RL-based methods instead optimize the policy gradient using reward signals, encouraging reasoning by exploring a much larger solution space. GRPO is closely related to PPO but differs in two key aspects: first, GRPO estimates the advantage using group-based estimation rather than a value function; second, it employs a set of fixed rules as the reward signal instead of a learned reward model. These optimizations make GRPO 50% more resource- and computation-efficient than PPO [30].
Definition.
@@
Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration. Let $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta_{\text{new}}}$ denote the old policy and current (new) policy, respectively, where $o$ is a complete response sampled from a policy. Let $\pi_{\theta_{\text{ref}}}$ denote the reference policy, which in practice is the frozen base MLLM. Let $G$ be the number of responses sampled per question in each iteration. The GRPO objective is given by:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (1)
@@
$\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\text{clip}\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\Bigg{)}A_{i}\Bigg{)}-\beta\mathbb{D}_{KL}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{)}$
@@
where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage. The KL divergence term [17] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$. Unlike PPO, which uses a critic model to estimate the advantage $A_{i}$ for a single response $o$, GRPO estimates the relative advantage by sampling a group of responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their rewards within the group to compute a relative advantage [30, 13]. Each reward is calculated based on rules without using a regard model. We detail the reward design below.
Reward design.
We follow [30] and use two types of reward: format and accuracy. To begin with, we prompt the model to explicitly output its thinking process in the “<think>…</think>" tag and the final answer in the “<answer>…</answer>" tag. The format reward is designed to check if the aforementioned tags are present in the final response. A reward score of 1 will be given if they exist and are correct. This helps the model to organize its thoughts and answer in a structured format for the ease of reading. The accuracy reward is a rule-based reward that checks if the actually answer matches with the ground truth. Similarly, a reward score of 1 is given when the results match. In practice, the ground truths are letter options “A, B, C, D" for multiple choice questions, and we treat all responses with correct letter options as the leading word as correct (“A…").
3.2 No-Thinking-Med-R1
Previous work [20] found that for the image classification task, compared with normal rule-based RL fine-tuning, excluding format accuracy and letting the model only output the final answer could lead to better performance. However, they only discussed this phenomenon for image classification task. In this paper, we further explore the thinking process of medical VQA and propose the No-Thinking-Med-R1.
@@ 3.2 No-Thinking-Med-R1
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs. The "Overall" column/row reports average performance per training and test domain.

### v4

@@ Abstract
Vision-language models (VLMs) have achieved impressive progress in natural image reasoning, yet their potential in medical imaging remains underexplored. Medical vision-language tasks demand precise understanding and clinically coherent answers, which are difficult to achieve due to the complexity of medical data and the scarcity of high-quality expert annotations. These challenges limit the effectiveness of conventional supervised fine-tuning (SFT) and Chain-of-Thought (CoT) strategies that work well in general domains. To address these challenges, we propose Med-R1, a reinforcement learning (RL)-enhanced vision-language model designed to improve generalization and reliability in medical reasoning. Built on the DeepSeek strategy, Med-R1 adopts Group Relative Policy Optimization (GRPO) to encourage reward-guided learning beyond static annotations. We comprehensively evaluate Med-R1 across eight distinct medical imaging modalities. Med-R1 achieves a 29.94% improvement in average accuracy over its base model Qwen2-VL-2B, and even outperforms Qwen2-VL-72B—a model with 36$\times$ more parameters. To assess cross-task generalization, we further evaluate Med-R1 on five question types. Med-R1 outperforms Qwen2-VL-2B by 32.06% in question-type generalization, also surpassing Qwen2-VL-72B. We further explore the thinking process in Med-R1, a crucial component for the success of Deepseek-R1. Our results show that omitting intermediate rationales (No-Thinking-Med-R1) not only improves in-domain and cross-domain generalization with less training, but also challenges the assumption that more reasoning always helps. These findings suggest that in medical VQA, it is not reasoning itself, but its quality and domain alignment, that determine effectiveness. Together, these results highlight that RL improves medical reasoning and generalization, enabling efficient and reliable VLMs for real-world deployment.
{IEEEkeywords}
Reinforcement Learning, Vision Language Models, Multimodal LLM, Post-Training, Medical Reasoning.
@@ 1 Introduction
[…]
Vision-language models (VLMs) have achieved strong performance on natural image understanding tasks such as visual question answering (VQA) and multimodal dialogue [1, 2, 3], enabled by large-scale pretraining and supervised fine-tuning (SFT). However, applying VLMs to medical imaging remains challenging due to the need for clinically sound interpretations and decision-making processes. Medical tasks often involve multi-step analysis—e.g., diagnosing a lung nodule may require integrating lesion localization, morphology, and context. In addition, the diversity of imaging modalities (e.g., CT, MRI) and task types (e.g., diagnosis, grading) imposes demands on generalizability. This raises a key question: how can we enable VLMs to perform well across medical domains while ensuring reliable and context-aware behavior?
@@
In this paper, we identify that the limitations of current medical VLMs primarily stem from the inherent drawbacks of Supervised Fine-Tuning (SFT) [4, 5]. While SFT has been widely adopted to adapt foundation models to medical imaging [6, 7, 8, 9, 10], it suffers from two fundamental issues that hinder medical applicability. First, SFT inherently biases models toward memorizing task-specific shortcuts rather than learning generalizable reasoning. By directly aligning model outputs with final answers (e.g., diagnostic labels), SFT encourages overfitting to superficial patterns in training data. Second, the scarcity of high-quality Chain-of-Thought (CoT) annotations severely limits the effectiveness of SFT in medical reasoning. Unlike general-domain tasks, where large-scale CoT datasets can be crowdsourced, medical reasoning requires domain-specific logical structuring (e.g., systematically ruling out differential diagnoses before confirming malignancy). However, curating such CoT datasets is prohibitively expensive, as it demands meticulous annotation by experienced medical professionals to ensure diagnostic validity and clinical coherence. As a result, existing SFT-based medical VLMs [6, 7] lack access to high-quality CoT data, leading to shallow reasoning with limited clinical rigor. These models frequently produce "black-box" predictions, struggling to provide traceable reasoning or maintain performance in out-of-domain tasks. This lack of transparency and robustness poses a significant challenge to medical adoption, where explainability and reliability are indispensable requirements.
To address these challenges, we propose Med-R1, a reinforcement learning (RL)-based framework for enhancing the generalizability and interpretability of medical VLMs. Unlike SFT, which aligns outputs to fixed supervision and often leads to shortcut learning, RL encourages exploration of diverse reasoning strategies through reward signals—without requiring explicit CoT annotations [11]. We adopt Group Relative Policy Optimization (GRPO) [12], a lightweight alternative to PPO [13] that stabilizes training via rule-based rewards and group-relative comparisons. These mechanisms reduce computational overhead while encouraging clinically grounded reasoning, making GRPO well-suited for medical tasks where scalability and reliability are essential.
@@
Figure 1: Overview of the Evaluation Framework: Eight Medical Imaging Modalities and Five Medical Vision Question Answering Tasks. We evaluate Med-R1 across eight distinct medical imaging modalities—CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy Images, and X-ray Imaging—as well as five medical vision question answering tasks: anatomy identification, disease diagnosis, lesion grading, modality recognition, and biological attribute analysis. Example images and corresponding clinical questions illustrate the diversity of medical data and the challenges in developing generalizable vision-language models for automated medical reasoning.
@@
As shown in Figure 1, we evaluate Med-R1 across eight diverse medical imaging modalities, including CT, MRI, Ultrasound, Dermoscopy, Fundus Photography, OCT, Microscopy, and X-ray. These modalities cover a wide range of clinical imaging—from macroscopic anatomy to cellular-level and functional assessments. Med-R1 (2B parameters) achieves 69.91% average accuracy, a 29.94% gain over its base model Qwen2-VL-2B, and even outperforms the 72B-parameter Qwen2-VL-72B (Table 3), highlighting the benefit of RL-driven adaptation. We further assess generalizability across five question types: modality recognition, anatomy identification, disease diagnosis, lesion grading, and biological attribute analysis. Med-R1 improves question-type generalization accuracy by 32.06% over Qwen2-VL-2B and surpasses Qwen2-VL-72B in this setting as well (Table 6). These demonstrate that RL enhances both parameter efficiency and task generalizability in medical VLMs.
Finally, inspired by recent findings on reasoning behavior in rule-based RL for classification tasks [14], we further investigate the role of intermediate rationales in Med-R1. A widely held assumption in VLM development is that generating step-by-step reasoning (“Think”) enhances generalization and performance. However, our empirical analysis reveals that this may not hold in domain-specific settings like medical VQA. We compare two RL post-training strategies: one that guides the model to produce explicit rationales before answering, and another that directly outputs the final answer without any intermediate reasoning steps. Surprisingly, we observe that the latter (No-Thinking) strategy often yields better generalization across both imaging modalities and question types. We hypothesize that in the absence of high-quality CoT supervision, RL-trained models fall back on their pre-trained reasoning patterns, which are typically tuned for general-domain tasks. Under large domain shifts—such as differences in image characteristics, task goals, and clinical language—this can result in hallucinated or misaligned rationales that impair performance. These findings suggest that it is not the presence of reasoning per se, but rather its quality and domain alignment, that determine its utility in medical vision-language reasoning.
@@
Med-R1 is a generalizable medical VLM that generates interpretable reasoning explanations for medical VQA across eight imaging modalities (Figure 2). We summarized our contribution as follows:
@@
    Multi-modality multi-task reasoning medical VLM. We propose Med-R1, the first reasoning medical VLM capable of supporting eight medical imaging modalities (CT, MRI, Ultrasound, etc.) across five unique tasks. We demonstrate that RL-based fine-tuning effectively promotes modality-specific as well as cross-modality reasoning in the medical domain without the need for token-level supervision as in SFT. The proposed Med-R1 is capable of generating step-by-step, accurate, and clinically plausible explanations.
@@
    Robust Generalization with Efficiency. We show that alongside the solid modality and task-specific performance, Med-R1 exhibits strong generalization. Med-R1 outperforms the base model by 29.94% and SFT baselines by 15.84% in average generalization accuracy across modalities. In cross-task settings, it outperforms the base model and SFT baseline by 32.06% and 11.25% respectively. Moreover, Med-R1 surpasses other larger generic or medical-specific models including Qwen2-VL-72B, and MedVInT-7B, warranting its efficiency and reliability for real-world deployment.
@@
    Rethinking the “More Thinking is Better” Assumption: Our results challenge the common belief that generating explicit thinking processes improves generalization. In medical VQA, we find that RL without reasoning steps yields more accurate predictions, while free-form reasoning—learned from general-domain data—can lead to hallucinations under domain shift. This highlights that in specialized domains—where image types, task goals, and language style differ significantly from general VLM training—the quality of reasoning, rather than its quantity, is critical for generalization.
@@ 2 Related Works
General VLMs and Medical VLMs. General-purpose VLMs such as CLIP [15] and BLIP-2 [16] have advanced natural image-text understanding via large-scale pretraining, but struggle with domain-specific tasks like medical reasoning. Recent efforts adapt VLMs through supervised fine-tuning (SFT) on medical datasets, as seen in LLaVA-Med [8] and Med-Flamingo [6]. While effective in-domain, these models often overfit to narrow corpora and lack generalization across modalities or task types. Our work addresses this limitation by introducing RL for scalable, modality-agnostic adaptation.
Reinforcement Learning for Post-Training. RL has shown promise for aligning language models with desired behavior via reward feedback [17, 18]. In vision-language tasks, RL improves VQA accuracy [11] and reduces hallucination [12], but often relies on complex reward models or costly human supervision. GRPO [12] offers a scalable alternative by using rule-based rewards and group-relative comparisons. We extend this approach to medical VQA, enabling efficient adaptation without modality-specific supervision.
Medical Reasoning and Interpretability Interpretable reasoning is critical in medical AI, with recent work exploring CoT prompting [19] and program-guided logic [4]. However, such methods rely on costly expert annotations [20], limiting scalability in clinical domains. Reinforcement learning offers an alternative by enabling emergent reasoning without explicit supervision. Concurrently, MedVLM-R1 [21] also applies GRPO-based RL to radiology VQA, which is conducted on a relatively small dataset without generalization analysis, making its reasoning robustness challenging to assess. Nonetheless, it highlights the growing interest in RL for medical VLMs.
Figure 2: Medical VQA examples of Med-R1 The left panel shows a modality recognition task, where the model correctly identifies the presence of breast tissue in a mammogram (X-ray). The right panel illustrates a disease diagnosis task, where Med-R1 accurately detects urolithiasis in an abdominal CT scan. In both cases, the model provides reasoning (“Think”), explaining its decision-making process before selecting the final answer.
@@ 3 Method
We adopt rule-based RL to encourage multimodal reasoning and generalization in medical domains. Recent studies have demonstrated that RL can incentivize emergent logical reasoning and generalization in multimodal tasks such as mathematical reasoning [22] and visual navigation [11]. Building on these insights, we extend its application to the medical domain and systematically assess its effectiveness in this context. Specifically, we leverage the popular RL-based post-training method GRPO [23] to train a large base MLLM [24] across 8 different modalities for medical reasoning and compare it with zero-shot and SFT performance of popular existing VLMs. We introduce the details of the algorithm, our reward design, and data structure below.
@@ 3.1 Group Relative Policy Optimization (GRPO)
Overview: RL-based algorithms such as PPO [13] and GRPO [12] belong to a family of fine-tuning and alignment strategies explicitly designed to enhance models’ reasoning capacities. Unlike supervised fine-tuning, which directly optimizes maximum likelihood, these RL-based methods instead optimize the policy gradient using reward signals, encouraging reasoning by exploring a much larger solution space. GRPO is closely related to PPO but differs in two key aspects: first, GRPO estimates the advantage using group-based estimation rather than a value function; second, it employs a set of fixed rules as the reward signal instead of a learned reward model. These optimizations make GRPO 50% more resource- and computation-efficient than PPO [12].
@@
Definition: Formally, let $P(Q)$ denote the question set used for training, where $q$ is a sampled question in the current iteration. Let $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta_{\text{new}}}$ denote the old policy and current (new) policy, respectively, where $o$ is a complete response sampled from a policy. Let $\pi_{\theta_{\text{ref}}}$ denote the reference policy, which in practice is the frozen base MLLM. Let $G$ be the number of responses sampled per question in each iteration. The GRPO objective is given by:
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\ \mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)}$
$\quad\frac{1}{G}\sum_{i=1}^{G}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},$
@@
$\qquad\text{clip}\left(\frac{\pi_{\theta_{\text{new}}}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)A_{i}\Bigg{)}$
$\quad\quad-\beta\,\mathbb{D}_{\text{KL}}\Big{(}\pi_{\theta_{\text{new}}}\Big{\|}\pi_{\text{ref}}\Big{)}\Bigg{]}$ (1)
@@
where $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the policy ratio, and $A_{i}$ is the estimated advantage, and $\epsilon$ is the clipping threshold for policy updates The KL divergence term [25] regularizes the policy update, ensuring that $\pi_{\theta}$ does not deviate excessively from the reference model $\pi_{\theta_{\text{ref}}}$. Unlike PPO, which uses a critic model to estimate the advantage $A_{i}$ for a single response $o$, GRPO estimates the relative advantage by sampling a group of responses $\{o_{i}\}_{i=1}^{G}$ and normalizing their rewards within the group to compute a relative advantage [12, 26]. Each reward is calculated based on rules without reward models. We detail reward design below.
@@
Table 1: Cross-Modality Generalization Performance of Med-R1 with RL Post-Training. Accuracy (%) across eight medical imaging modalities, where rows indicate training modalities and columns test modalities. Darker cell shades indicate higher accuracy for corresponding training-test pairs in each column. The base model is Qwen2-VL-2B.